OpenStudy (anonymous):

Jason has 50 coins, all in 5 and 1 peso coins, amounting 130 php. How many 5 peso coins does he have?

OpenStudy (amistre64):

what is 130 php? a total value?

OpenStudy (anonymous):

I have no formula for it but I think he has 20 Php 5 peso coins :)

OpenStudy (amistre64):

p @ 5 = (130 - 50(1))/(5-1) is my guess

OpenStudy (amistre64):

what is a "php" ?

OpenStudy (anonymous):

Because 20*5= 100 Then he has 30 Php 1 coins Then 30*1= 30 So 100+30= 130. :)

OpenStudy (anonymous):

Php= Philippine Currency

OpenStudy (amistre64):

is a peso a phillipian coin? cause it sounds mexicoian to me

OpenStudy (amistre64):

what I get is 20 pesos as well

OpenStudy (anonymous):

Yeah, its Philippine.

OpenStudy (amistre64):

the long route is to set up a system of equations such that: a + b = t ; fives + ones = 50 ax + by = z ; x and y are the number of each and z = 130, now eliminate to find your a and b -ax -bx = -tx ax +by = z -------------- b(y-x) = z - tx b@y = (z-tx)/(y-x) p @ 5 = (130 - 50(1))/(5-1) = 20 pesos @ a 5 count

OpenStudy (amistre64):

that needs an edit; x and y are the values for each coin; 5 and 1

OpenStudy (anonymous):

What amount should an employee receive as bonus so that he would net 10 000 pesos after deducting 30 % for taxes?

OpenStudy (amistre64):

10000/1.30 = original amount

OpenStudy (amistre64):

A = P(1+r) ; P = A/(1+r)

OpenStudy (anonymous):

solutiom please

OpenStudy (amistre64):

im sure that you can use a calculator at least; i have faith in you :)

OpenStudy (anonymous):

yes.. what is the ssolution